OpenStudy (anonymous):

solve the equation, tan x = 1, for the given interval of x: -pi/2 < x < pi/2

OpenStudy (anonymous):

\[-\frac{ \pi }{ 2 } \le x \le \frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

just to clarify this is what the problem really looks like in equation form.

OpenStudy (anonymous):

Remember the special angles :D To find x, use arc tan 1 =x

OpenStudy (anonymous):

I think i figured it to be pi/4... is that correct?

OpenStudy (anonymous):

Yea pi/4 is correct but you have to make are there any other answers

OpenStudy (anonymous):

I couldn't come up with anything is, I believe this is the only solution

OpenStudy (anonymous):

anything else*

OpenStudy (anonymous):

Look at the given interval, are there any other solution which gives you a 1

OpenStudy (anonymous):

oops sorry, i saw intervals wrongly, so yea there's only one solution